If three times the first number plus four times the second number equals 26, and ten times the first number and minus two times the second number equals 10, what are the numbers?: If three times the first number plus four times the second number equals 26, and ten times the first number and minus two times the second number equals 10, what are the numbers?
Answer by ankor@dixie-net.com(4538)

Let x = 1st number
Let y = 2nd number
:
Write an equation for each statement:
"If three times the first number plus four times the second number equals 26,"
3x + 4y = 26
:
"ten times the first number and minus two times the second number equals 10,"
10x - 2y = 10
:
what are the numbers?
;
We can use elimination here; multiply the 2nd equation by 2 and add to the 1st
20x - 4y = 20
3x + 4y = 26
-----------------adding eliminates y
23x + 0y = 46
x = 46/23
x = 2
:
Use the 1st equation to find y, (substitute 2 for x), Check the solutions in the 2nd equation.